a cook cuts a cake into 35 equal pieces and an identical cake into 42 pieces. The caterer, however, insists that the cakes be cut exactly alike. Into how many pieces must each cake now be cut
Answer by khwang(438)

To find the LCM of 35 and 42
35 = 5*7, 42 = 2*3*7,
the LCM is 2*3*5*7 = 210
More precisely, if we cut each piece (1/35) of the 1st cake by 210/35
= 6 parts equally and cut each piece (1/42) of the 2nd cake by 210/42
= 5 partse qually. Then the two cakes will be cut exactly alike.
(each with 210 equal pieces)
Kenny